OpenStudy (anonymous):

3 3/4 divided by 1 1/2 =

OpenStudy (anonymous):

wow i just figured it out

OpenStudy (anonymous):

3+ 3/4 = 15/4 1 + 1/2 = 3/2 (15/4)/(3/2) = 5/2

OpenStudy (anonymous):

haha nevermind i was wrong

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

No problem :)

OpenStudy (anonymous):

Can you help me with another?

OpenStudy (anonymous):

Sure